# 3D Rigid Body Localization and Motion Estimation via Gaussian Belief Propagation

### v1

@@ Abstract
We propose a novel solution to the moving rigid body localization (RBL) problem, in which the three-dimensional (3D) rotation angles and the translation distances, as well as the angular and translational velocities are estimated by only utilizing the relative range and Doppler measurements between the “anchor” sensors at the observer and “target” sensors of the rigid body. The proposed method is based on a bilinear Gaussian belief propagation (GaBP) framework, employed to estimate the absolute sensor positions and velocities using a range- and Doppler-based received signal model, which is then utilized in the reconstruction of the RBL transformation model, linearized under a small-angle approximation. The method further incorporates a second bivariate GaBP designed to directly estimate the 3D rotation angles and translation distances, including an interference cancellation (IC) refinement stage to improve the angle estimation performance, followed by the estimation of the angular and the translational velocities. The effectiveness of the proposed method is verified via simulations, which confirms its improved performance compared to equivalent state-of-the-art (SotA) techniques.
@@
Rigid body localization (RBL), Motion estimation, Wireless sensor network (WSN), Range-Doppler-based positioning, Gaussian belief propagation (GaBP).
@@ I Introduction
Recent years have seen great advancements in wireless sensor technologies, which are capable of detecting environmental parameters such as temperature, pressure, luminosity, humidity and the strength of electric signals, and directly transmitting them to one or multiple receivers comprising a wireless sensor network (WSN), for applications ranging from monitoring and control, to smart factories, to Internet-of-Things (IoT), to positioning systems [2, 3, 4] and more. In fact, many WSN applications either require or can be improved under the knowledge of accurate location information, such that the sensor location problem has been studied extensively [5, 6].
@@
More recent sensing-related applications such as virtual reality (VR), extended reality (XR), robotics and autonomous vehicles, however, require not only precise location information of individual sensors, but also the orientation of the various sensors associated to a given body or object [7, 8], giving rise to a variation of the positioning problem known as rigid body localization (RBL) [9, 10, 11, 12, 13, 14], where the relative position of sets of sensors is fixed according to a rigid conformation, whose translation and rotation (i.e., orientation) must be estimated.
@@
Several effective strategies exists to estimate the location and orientation of objects, including computer vision-based techniques, which focus on feature extraction and posture estimation [15, 16], usually based on image/video signals; and inertial measurement unit (IMU)-based techniques, which leverage information from accelerometers, gyroscopes, and magnetometers [17, 18]. A problem of computer-vision approaches is, however, that they typically require high volumes of data and rely on high-complexity methods, which limits their application. In turn, IMU-based methods are often unreliable or inaccurate, requiring frequent sensor re-calibrations, not to mention the aid of external radio technology, which to some extent defeats the self-reliant idea behind the approach.
In contrast to the latter, the RBL concept considered here [9, 10, 11, 12, 13, 14] aims to generalize the traditional localization paradigm and exploits range measurements between sensors on the rigid body and a set of anchor sensors at known positions [19, 20], to estimate the location, orientation (and possibly the shape) of rigid bodies.
@@
Earlier state-of-the-art (SotA) RBL is based on either algebraic methods leveraging multidimensional scaling (MDS) [21] or semidefinite relaxation (SDR) [22, 23], both of which have a cubic computational complexity at a minumum. Although least squares methods such as those in [24, 25, 26] offer lower complexity alternatives to solve the RBL problem, all the aforementioned approaches have in common the fact that they only take into account stationary scenarios, in the sense that they seek to estimate only the translation and the orientation of the rigid body, respectively defined as the distance and the rotation matrix of the rigid body relative to a given reference.
@@
In contrast, in RBL schemes suitable to moving scenarios, one must in addition to translation and orientation, also estimate the angular and translational velocities. Unfortunately, existing moving RBL solutions are either of low complexity but low precision, as is the case of the alternating minimization-based method proposed in [27], or of high precision but also of high complexity, such as the SDR method of [28].
@@
A rare example of a well-balanced solution – $i.e.$, namely, with both high precision and low complexity – for both stationary and moving RBL is, however, the method proposed in [11], where preliminary position estimates for each sensor is first obtained from range measurements using a least-squares formulation based on the linearized model described in [29], followed by the extraction of rotation and translation parameters via singular value decomposition (SVD)-based analysis and finally a conformation-enforcing refinement based on an Euler angles formulation, solved via a weighted least squares (WLS) approach. However, this stationary RBL technique is further extended into a two stage solution for the moving RBL scenario, in which the preliminary range-based position estimation stage is upgraded to include also velocity estimates for each sensor obtained from Doppler measurements, followed by the estimation of angular and translational velocities using the linearized model described in [30].
@@
In light of the above, we propose a series of novel algorithms towards solving stationary and moving RBL problems in this article. First, two methods for the location and velocity estimation of sensors in three-dimensional (3D) space are produced. In particular, the techniques, labeled Algorithms 1 and 2, respectively, are designed for the estimation of 3D rotation angles and translation distances from range measurements, as well as angular and translational velocities from Doppler measurements. Then, two bivariate linear Gaussian belief propagation (GaBP) algorithms are offered for the stationary and moving RBL problems. Those contributions, dubbed Algorithms 4 and 3, respectively, are capable of estimating the 3D rotation angles and translation distances, and respectively the 3D angular and translational velocities, utilizing range/Doppler measurements and the estimated sensor position and velocities from Algorithms 1 and 2 by leveraging bilinear GaBP over a small-angle modification of the linearized system models of [30] and [29], with the rigid body conformation constraint directly incorporated. The bivariate methods in Algorithms 4 and 3, respectively, are shown to outperform the SotA two-stage RBL methods in terms of parameter estimation performance, while retaining low computational complexity.
@@ III Proposed Rigid Body Localization
In this section, we propose a low-complexity position and transformation estimator for RBL, in light of the system model derived in Section II and by leveraging the GaBP message passing framework. Specifically, an initial GaBP is derived to solve the position-explicit system in equation (LABEL:eq:linear_sys) to obtain sensor coordinates. Then, in possession of the initial position estimate, a second GaBP is derived on the rigid body transformation parameter-explicit system in equation (28), to obtain the final estimate of the 3D rotation angles $\bm{\theta}$ and translation distances $\bm{t}$.
@@ V-A RBL Parameter Estimation Performance Against the SotA
Firstly, for the stationary scenario, Figure 3 illustrates the sensor position estimation performance of the proposed Algorithm 1 and the SotA closed-form solution of [29] for different noise standard deviations, also referred to as range error in meters. It can be observed that the proposed linear GaBP method outperforms the SotA approach for all range error regimes, which suggests that already the proposed preliminary positioning method can be used as an initializer for other SotA methods to perform RBL parameter extraction. In addition, Figures 5 and 5 illustrate the estimation performance of the rigid body translation distances and rotation angles respectively.
@@ V-B Convergence Behavior and Complexity Analysis
Here, the performance of the proposed RBL parameter estimation algorithms are further assessed in terms of the computational complexity and convergence behavior, and compared against the SotA methods.
@@ VI Conclusion
We presented a novel and efficient framework for solving the RBL problem, for stationary and moving rigid bodies, i.e., estimation of position, 3D rotation, translation, sensor velocity, angular velocity and translational velocity via a series of tailored GaBP message passing estimators. First, a linear GaBP for the sensor position estimation based on the obtained range estimates from the wireless sensors is derived, from which the RBL system is reformulated via the small-angle approximation to enable the construction of a second bivariate GaBP which is capable of directly estimating the rotation angles and the translation distances. Next, Doppler measurements were obtained, designing a system of linear equations that can be used to estimate the sensor velocity via a linear GaBP, followed by the estimation of angular and translational velocity by the construction of another bivariate GaBP. The proposed RBL method is shown to outperform the SotA method in all of the position, rotation, translation, sensor velocity, angular velocity and translational velocity estimation performance, in addition to the complexity advantage highlighted by the analysis.

### v2

# 6D Rigid Body Localization and Velocity Estimation via Gaussian Belief Propagation

@@ Abstract
We propose a novel message-passing solution to the sixth-dimensional (6D) moving rigid body localization (RBL) problem, in which the three-dimensional (3D) translation vector and rotation angles, as well as their corresponding translational and angular velocities, are all estimated by only utilizing the relative range and Doppler measurements between the “anchor” sensors located at an 3D (rigid body) observer and the “target” sensors of another rigid body. The proposed method is based on a bilinear Gaussian belief propagation (GaBP) framework, employed to estimate the absolute sensor positions and velocities using a range- and Doppler-based received signal model, which is then utilized in the reconstruction of the RBL transformation model, linearized under a small-angle approximation. The method further incorporates a second bivariate GaBP designed to directly estimate the 3D rotation angles and translation vectors, including an interference cancellation (IC) refinement stage to improve the angle estimation performance, followed by the estimation of the angular and the translational velocities. The effectiveness of the proposed method is verified via simulations, which confirms its improved performance compared to equivalent state-of-the-art (SotA) techniques.
@@
Rigid body localization (RBL), Message Passing, 6D Localization, Attitude Estimation.
@@ I Introduction
Recent years have seen great advancements in wireless sensor technologies, which are capable of detecting environmental parameters such as temperature, pressure, luminosity, humidity and the strength of electric signals, that are used in applications ranging from monitoring and control, to smart factories, to Internet-of-Things (IoT), to positioning systems [2, 3, 4] and more. In fact, many wireless sensor network (WSN) applications either require or can be improved under the knowledge of accurate location information, such that the sensor location problem has been studied extensively [5, 6].
@@
More recent sensing-related applications such as virtual reality (VR), extended reality (XR), robotics and autonomous vehicles, however, require not only precise location information of individual sensors, but also the orientation of the various sensors associated to a given body or object [7, 8], giving rise to a variation of the positioning problem known as rigid body localization (RBL) [9, 10, 11, 12, 13, 14], where the relative position of sets of sensors is fixed according to a rigid conformation, whose translation and rotation (i.e., orientation or attitude) must be estimated.
@@
Several effective strategies exists to estimate the location and attitude of objects, including computer vision-based techniques, which focus on feature extraction and posture estimation [15, 16], usually based on image/video signals; and inertial measurement unit (IMU)-based techniques, which leverage information from accelerometers, gyroscopes, and magnetometers [17, 18]. A problem of computer-vision approaches is, however, that they typically require high volumes of data and rely on high-complexity methods, which limits their application. In turn, IMU-based methods are often unreliable or inaccurate, requiring frequent sensor re-calibrations, not to mention the aid of external radio technology, which to some extent defeats the self-reliant idea behind the approach.
@@
In contrast to the latter, the RBL concept considered here [9, 10, 11, 12, 13, 14] aims to generalize the traditional localization paradigm, and exploits range measurements between sensors on the rigid body and a set of anchor sensors at known positions [19, 20], to estimate the location, orientation (and possibly the shape) of rigid bodies. Earlier state-of-the-art (SotA) contributions on such RBL approach are, however, based on either algebraic methods leveraging multidimensional scaling (MDS) [21] or semidefinite relaxation (SDR) [22, 23], both of which have a cubic computational complexity at least. And although least squares methods [24, 25, 26] have been used to lower the complexity of the RBL problem, these techniques have in common the fact that they only take into account stationary scenarios, in the sense that they seek to estimate only the translation and the orientation ($i.e.$, the pose) of the rigid body, respectively defined as the distance and the rotation matrix of the rigid body relative to a given reference.
@@
In contrast, in RBL schemes suitable to moving scenarios one must, in addition to pose, also estimate the corresponding angular and translational velocities. Unfortunately, existing moving RBL solutions are either of low complexity but low precision, as is the case of the alternating minimization-based method proposed in [27], or of high precision but also of high complexity, such as the SDR method of [28].
@@
An example of a well-balanced solution – $i.e.$, namely, with both high precision and low complexity – for both stationary and moving RBL is, however, the method proposed in [11], where preliminary position estimates for each sensor is first obtained from range measurements using a least-squares formulation based on the linearized model described in [29], followed by the extraction of rotation and translation parameters via singular value decomposition (SVD)-based analysis and finally a conformation-enforcing refinement based on an Euler angles formulation, solved via a weighted least squares (WLS) approach. However, this stationary RBL technique is further extended into a two stage solution for the moving RBL scenario, in which the preliminary range-based position estimation stage is upgraded to include also velocity estimates for each sensor obtained from Doppler measurements, followed by the estimation of angular and translational velocities using the linearized model described in [30].
@@
In light of the above, and inspired by former message passing based localization approaches [37, 38, 39], we propose here a series of novel algorithms to solve the stationary and moving RBL problems described in the article. First, two methods for the location and velocity estimation of sensors in three-dimensional (3D) space are produced. In particular, the techniques, labeled Algorithms 1 and 3, respectively, are designed for the estimation of sensor positions from range measurements, as well as sensor velocities from Doppler measurements. Then, two bivariate linear Gaussian belief propagation (GaBP) algorithms are offered for the stationary and moving RBL problems. Those contributions, dubbed Algorithms 2 and 4, respectively, are capable of estimating the 3D rotation angles and translation distances, and respectively the 3D angular and translational velocities, utilizing range/Doppler measurements and the estimated sensor position and velocities from Algorithms 1 and 2 by leveraging bilinear GaBP over a small-angle modification of the linearized system models of [30] and [29], with the rigid body conformation constraint directly incorporated. The bivariate methods in Algorithms 4 and 3, respectively, are shown to outperform the SotA two-stage RBL methods in terms of parameter estimation performance, while retaining low computational complexity.
@@ III Proposed Rigid Body Localization
In this section, we propose a low-complexity position and transformation estimator for RBL, in light of the system model derived in Section II and by leveraging the GaBP message passing framework. Specifically, an initial GaBP is derived to solve the position-explicit system in equation (LABEL:eq:linear_sys) to obtain sensor coordinates. Then, in possession of the initial position estimate, a second GaBP is derived on the rigid body transformation parameter-explicit system in equation (28), to obtain the final estimate of the 3D rotation angles $\bm{\theta}$ and translation vector $\bm{t}$.
@@ V-A RBL Parameter Estimation Performance Against the SotA
Firstly, for the stationary scenario, Figure 3 illustrates the sensor position estimation performance of the proposed Algorithm 1 and the SotA closed-form solution of [29] for different noise standard deviations, also referred to as range error in meters. It can be observed that the proposed linear GaBP method outperforms the SotA approach for all range error regimes, which suggests that already the proposed preliminary positioning method can be used as an initializer for other SotA methods to perform RBL parameter extraction. In addition, Figures 5 and 5 illustrate the estimation performance of the rigid body translation vector and rotation angles respectively.
@@ V-B Convergence Behavior and Complexity Analysis
Here, the performance of the proposed RBL parameter estimation algorithms are further assessed in terms of the computational complexity and convergence behavior, which is thoroughly investigated in [42, 43, 44], and compared against the SotA methods.
@@ VI Conclusion
We presented a novel and efficient framework for solving the sixth-dimensional (6D) RBL problem, for under stationary and moving i.e., estimation of 3D location and 3D rotation, as well as corresponding translation and angular velocities, via a series of tailored GaBP message passing estimators. First, a linear GaBP for the sensor position estimation based on the obtained range estimates from the wireless sensors is derived, from which the RBL system is reformulated via the small-angle approximation to enable the construction of a second bivariate GaBP which is capable of directly estimating the rotation angles and the translation distances. Next, Doppler measurements were obtained, designing a system of linear equations that can be used to estimate the sensor velocity via a linear GaBP, followed by the estimation of angular and translational velocity by the construction of another bivariate GaBP. The proposed RBL method is shown to outperform the SotA method in all of the position, rotation, translation, sensor velocity, angular velocity and translational velocity estimation performance, in addition to the complexity advantage highlighted by the analysis.